OpenStudy (anonymous):

solve for x in the given interval sinxtan3x=0(-pi,0)

OpenStudy (anonymous):

given sinx tan3x=0 means either sinx=0 or tan3x=0 if sinx=0 then x=n(pi) where n is an integer if tan3x=0 then 3x=m(pi) where m is an integer for getting the values in the given inter use proper values for n and m and thus the answer are 0.-(pi)/3,-2(pi)/3,-pi etc..

OpenStudy (anonymous):

how is -2pi/3 in the interval?

OpenStudy (anonymous):

when m=2 in 3x=mpi and it is less than -pi also tan3(-2(pi)/3))=0